OpenStudy (anonymous):

Turn 15x+5y=-30 into slope-intercept form.

OpenStudy (zzr0ck3r):

Subtract \(15x\) from both sides and tell me what you have.

OpenStudy (anonymous):

5y=15x+30?

OpenStudy (zzr0ck3r):

subtract

OpenStudy (zzr0ck3r):

why did you change 30 to -30?

OpenStudy (anonymous):

sorry.. so is it 5y=15x+-30

OpenStudy (zzr0ck3r):

\(15x+5y=-30 \\ 15x+5y-15x=-30-15x\\5y=-30-15x\) Are you with me?

OpenStudy (anonymous):

Yes!

OpenStudy (zzr0ck3r):

Now divide every term by \(5\).

OpenStudy (anonymous):

okay so y=-6x-3

OpenStudy (zzr0ck3r):

you changed the term that the x was in

OpenStudy (anonymous):

.............??

OpenStudy (zzr0ck3r):

what is -30 divided by 5?

OpenStudy (anonymous):

oh... -6

OpenStudy (zzr0ck3r):

Yes, not -6x

OpenStudy (zzr0ck3r):

but -15x divided by 5 is -3x

OpenStudy (zzr0ck3r):

so what do we have?

OpenStudy (anonymous):

-3x+-6

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

THANKS

OpenStudy (zzr0ck3r):

np